# ToonifyGB: StyleGAN-based Gaussian Blendshapes for 3D Stylized Head Avatars

### v1

@@ Abstract.
The introduction of 3D Gaussian blendshapes has enabled the real-time reconstruction of animatable head avatars from monocular video. Toonify, a StyleGAN-based framework, has become widely used for facial image stylization. To extend Toonify for synthesizing diverse stylized 3D head avatars using Gaussian blendshapes, we propose an efficient two-stage framework, ToonifyGB. In Stage 1 (stylized video generation), we employ an improved StyleGAN to generate the stylized video from the input video frames, which addresses the limitation of cropping aligned faces at a fixed resolution as preprocessing for normal StyleGAN. This process provides a more stable video, which enables Gaussian blendshapes to better capture the high-frequency details of the video frames, and efficiently generate high-quality animation in the next stage. In Stage 2 (Gaussian blendshapes synthesis), we learn a stylized neutral head model and a set of expression blendshapes from the generated video. By combining the neutral head model with expression blendshapes, ToonifyGB can efficiently render stylized avatars with arbitrary expressions. We validate the effectiveness of ToonifyGB on the benchmark dataset using two styles: Arcane and Pixar.
Toonify, StyleGAN, Gaussian Splatting, Gaussian Blendshapes, 3D Stylized Head Avatar, Facial Animation
journal: TOG
1. Introduction
@@
With the advancement of 3D head reconstruction technologies, individuals are now able to personalize unique avatars for telepresence and virtual/augmented reality (VR/AR) applications, which serve as an essential foundation for the rise of the metaverse. Considering user preferences and privacy concerns, the creation of stylized avatars is an important topic that merits further investigation. Toonify (Pinkney and Adler, 2020) is a StyleGAN‑based method for stylizing 2D facial images. In contrast to photo-realistic 3D head avatars, stylized 3D avatars emphasize the expression of personal identity and the faithful transfer of target styles.
@@
Blendshape is an efficient facial animation representation method that can generate continuous and high-quality expressions through a set of 3D meshes, each corresponding to a specific facial expression. These facial shapes are generated by linearly blending the basis meshes using weighting coefficients. With the advent of Neural Radiance Fields (NeRF) (Mildenhall et al., 2021), Gao et al. (Gao et al., 2022) and Zheng et al. (Zheng et al., 2022) introduced the blendshape concept into NeRF, enabling avatar animation through the construction of a group of NeRF blendshapes that are linearly blended. Furthermore, the recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) significantly improves rendering efficiency while achieving high-quality head reconstruction compared to NeRF. Based on this, 3D Gaussian Blendshapes (3DGB) (Ma et al., 2024) successfully integrated the blendshape concept with Gaussian splatting, achieving both fast rendering speeds and state-of-the-art performance in head reconstruction and animation.
@@
In contrast to the aforementioned studies that focus on reconstructing photo-realistic head avatars, we propose ToonifyGB, a two-stage framework for stylized 3D head avatar synthesis and animation. Given monocular input video frames, Stage 1 employs an improved StyleGAN to generate a more stable (i.e., less jittery) stylized video beyond pre‑aligned cropping. In Stage 2, we begin with 3DGB to reconstruct a neutral head model with a set of expression blendshapes, each represented as 3D Gaussians, as shown in 1. Finally, by incorporating the facial tracker (Zielonka et al., 2022), we use the tracked motion parameters to drive ToonifyGB in synthesizing the animated head avatar.
@@
  • We propose ToonifyGB, an efficient two-stage framework for synthesizing 3D stylized head avatars from input videos using Gaussian blendshapes. This framework supports diverse toonified styles and enables real-time animation.
@@
  • We demonstrate that reducing frame-by-frame jitter enables Gaussian blendshapes to better capture the high-frequency details in the generated video, improving the quality of 3D stylized head avatar animations.
@@
  • To the best of our knowledge, we are the first to synthesize 3D stylized head avatars using Gaussian blendshapes.
2. Related Work
2.1. StyleGAN and Toonify
StyleGAN (Karras et al., 2019, 2020) has been widely used to generate realistic facial images across diverse styles. Inversion of StyleGAN enables projecting real facial images into its latent space, allowing for subsequent edits such as adding glasses or changing hairstyle or age (Abdal et al., 2019; Patashnik et al., 2021). To enhance inversion efficiency, methods such as pSp (Richardson et al., 2021) and e4e (Tov et al., 2021) employ encoders to directly project target faces into their corresponding latent codes. However, these methods often struggle to reconstruct fine image details, resulting in unsatisfactory reconstruction quality. To address these limitations, ReStyle (Alaluf et al., 2021) and HFGI (Wang et al., 2022) improve reconstruction fidelity by respectively predicting latent code residuals and correcting middle layer features. Nevertheless, these methods remain limited to aligned and cropped facial images to achieve effective editing and reconstruction.
Recently, researchers (Pinkney and Adler, 2020; Ojha et al., 2021; Jang et al., 2021; Yang et al., 2022a; Gal et al., 2022) have explored using StyleGAN for target‑domain image generation through transfer learning. Among these works, Toonify (Pinkney and Adler, 2020) fine‑tunes the trained generator to blend realistic textures with toonified facial structures. In addition to image editing, StyleGAN has also been widely applied to video editing. Related studies have focused on enhancing video editing performance by employing temporal correlations in low‑dimensional latent codes (Fox et al., 2021), disentangling identity from facial attributes (Yao et al., 2021), incorporating sketch‑based branches (Liu et al., 2022), and tuning the generator to maintain temporal consistency (Tzaban et al., 2022). However, these methods typically rely on facial alignment and cropping as preprocessing steps. Although StyleGAN3 (Karras et al., 2021) was introduced to support unaligned face inputs, a subsequent study (Alaluf et al., 2022) has shown that it struggles to encode facial features effectively without preprocessing, often resulting in structural artifacts. To overcome these limitations, methods such as VToonify (Yang et al., 2022b) and StyleGANEX (Yang et al., 2023) have been proposed to directly process videos beyond pre-aligned cropping. Nevertheless, these methods remain limited to 2D image representations and have yet to be extended to 3D applications.
@@ 2.1. StyleGAN and Toonify
Figure 2. Pipeline: Our ToonifyGB framework consists of two stages: Stage 1 involves the generation of stylized videos, and Stage 2 focuses on the synthesis of 3D stylized head avatars using Gaussian blendshapes.
[…]
The architecture of our method ToonifyGB.
2.2. 3D Head Avatar
@@
Since the introduction of NeRF (Mildenhall et al., 2021), implicit representation-based methods (Yenamandra et al., 2021; Zheng et al., 2022; Hong et al., 2022; Chan et al., 2022; Xu et al., 2023) for head reconstruction have achieved remarkable progress. 3DGS (Kerbl et al., 2023) has obtained a significant breakthrough in 3D reconstruction, further advancing the development of downstream applications such as 3D head modeling. Although several Gaussian-based methods (Qian et al., 2024; Xu et al., 2024; Chen et al., 2024; Ma et al., 2024; Xiang et al., 2024; Abdal et al., 2024; Kirschstein et al., 2024) have demonstrated high-quality head reconstruction and impressive rendering performance, they typically focus on photo-realistic avatars, with relatively limited exploration of avatar stylization. Stylized head avatars, characterized by geometric abstraction and artistic expression, differ significantly from the photo-realistic avatars synthesized by the aforementioned methods.
@@
Pre-trained 3D GANs (Wu et al., 2016) enable high-quality generation, making 3D head stylization possible. Although fine-tuning 3D generators for geometric and texture-based stylization has proven effective (Or-El et al., 2022; Jin et al., 2022; Abdal et al., 2023; Lan et al., 2023), performing independent fine-tuning for each new style remains costly. Toonify3D (Jang et al., 2024) addressed this limitation by predicting facial surface normals using the proposed StyleNormal, enabling direct face stylization without additional fine-tuning. Similarly, DeformToon3D (Zhang et al., 2023) introduced StyleField to predict conditional 3D deformations, aligning NeRF representations in real space with style space to achieve geometric stylization and obviate per-style fine‑tuning. However, Toonify3D suffers from limited data diversity, and DeformToon3D cannot support novel-view animations, which limits their application scenarios.
3. Method
3.1. ToonifyGB Framework
@@
Given a monocular video input, ToonifyGB applies frame-by-frame stylization to generate the corresponding stylized frames. For inputs such as live streams or selfie videos, the face often occupies only a small portion of each frame, while the rest includes the hairstyle and upper body. Traditional methods (Karras et al., 2019, 2020) typically require facial alignment, cropping, and editing before synthesizing the results back into the original frame. This process often introduces visual discontinuities at the seams, resulting in noticeable jitter in the output video. To address this problem, we employ an improved StyleGAN based on StyleGANEX (Yang et al., 2023) in Stage 1, enabling stable stylized video generation at the original resolution, as shown in Figure 2.
@@
To prepare the training data for Stage 2, we follow (Zielonka et al., 2023; Ma et al., 2024), using the facial tracker from (Zielonka et al., 2022) to compute FLAME (Li et al., 2017) meshes for a neutral head model and a set of basis expression blendshapes. This process also provides camera parameters $C$, joint and pose parameters $\Theta$, and expression coefficients $\{\psi_{k}\}$ for each frame. In addition to enabling facial expressions control, the FLAME model based on Principal Component Analysis (PCA) provides joint and pose parameters for controlling head, eyeball, eyelid, and jaw movements. As shown in Figure 2, we apply Linear Blend Skinning (LBS) to transform 3D head avatar: $B^{\psi*}=LBS(B^{\psi},\Theta)$. The transformed Gaussian model is then rendered into a Gaussian head avatar in real-time using Gaussian Splatting. Finally, by incorporating camera parameters, we enable head avatar animation from novel views.
Figure 3. Generators of StyleGAN2 and ours: For clarity and ease of comparison, we omit components that are exactly the same: the learned weights, biases, and noises.
[…]
StyleGAN2 generator vs. our generator.
3.2. Stylized Video Generation
@@
As demonstrated by StyleGANEX (Yang et al., 2023), the fixed $4\times 4$ resolution of the first‑layer feature map in StyleGAN2 (Karras et al., 2020) is insufficient for capturing fine details in non‑aligned faces. In contrast, the seventh layer produces a higher‑resolution $32\times 32$ feature map, enabling more accurate facial generation. Based on this, we propose enhancing the feature maps of the seven shallow layers in StyleGAN2 to the $32\times 32$ resolution. As shown in Figure 3, our improvements to StyleGAN2 are as follows. First, we replace the constant $4\times 4$ input of the first layer with a variable feature map of resolution $H/32\times W/32$, thus supporting arbitrary input sizes. Then, we substitute the standard convolutions in the shallow layers with dilated convolutions to enlarge the receptive field. Finally, we remove all upsample operations before the eighth layer, ensuring that the seven shallow layers maintain the same $32\times 32$ resolution. These improvements effectively address the limitations beyond pre-aligned cropping. Our improved StyleGAN can stylize input videos without requiring facial alignment and cropping as preprocessing, as shown in Figure 9.
3.3. Gaussian Blendshapes Synthesis
@@
We represent all Gaussian head avatars using 3D Gaussians. As described in 3DGS (Kerbl et al., 2023), each Gaussian has some basic properties including Gaussian center $\mu$, scale $s$, color $c$, opacity $\alpha$, and rotation $q$. Based on 3DGB (Ma et al., 2024), our Gaussian blendshape representation consists of a neutral model $B_{0}$ and a set of $n$ expression blendshapes ${B_{1},B_{2},\ldots,B_{n}}$. Each Gaussian in the neutral model $B_{0}$ has a set of blend weights $w$ to control joint and pose. In addition, each Gaussian in an expression blendshape $B_{k}$ corresponds one‑to‑one to a Gaussian in the neutral model $B_{0}$. The difference between $B_{k}$ and $B_{0}$ is defined as the difference in their corresponding Gaussian properties: $\Delta B_{k}=B_{k}-B_{0}$. The expression of Gaussian head avatar $B^{\psi}$ can be computed as follows:
(1) $B^{\psi}=B_{0}+\sum_{k=1}^{n}\psi_{k}\,\Delta B_{k}$
where $\psi_{k}$ are the expression coefficients.
@@
Since the FLAME meshes and blendshape models do not include interior mouth components such as teeth, we adopt the method of 3DGB (Ma et al., 2024) by defining a separate set of Gaussians for the mouth $B_{m}$. The properties of these mouth Gaussians are not affected by expression changes, they only move with the jaw joint in the FLAME model. The mouth for the Gaussian head avatar $B_{m}^{*}$ is computed as:
(2) $B_{m}^{*}=\mathrm{LBS}(B_{m},\Theta),$
where $\mathrm{LBS}$ denotes linear blend skinning with the joint and pose parameters $\Theta$. In summary, both $B^{\psi*}$ and $B_{m}^{*}$ are rendered together to form the complete Gaussian head avatar, as shown in Figure 2.
3.4. Loss Function
We adopt the loss function from 3DGB (Ma et al., 2024), and define the total loss as follows:
(3) $L=\lambda_{1}L_{rgb}+\lambda_{2}L_{\alpha}+\lambda_{3}L_{reg},$
where the default values of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are 1, 10, 100, respectively.
@@ 3.4. Loss Function
$L_{rgb}$ encourages the rendered image to match the video frame in both color and structure. It is computed as a weighted blend of an $L_{1}$ loss and a D-SSIM loss:
(4) $L_{rgb}=\lambda_{rgb}L_{1}+(1-\lambda_{rgb})L_{D-SSIM},$
where the default value of $\lambda_{rgb}$ is 0.2.
@@
$L_{\alpha}$ penalizes opacity outside the head mask. For each frame $i$, we compute the accumulated opacity image $I_{\alpha}$ and head mask $M_{h}^{i}$, then average over $F$ frames:
(5) $L_{\alpha}=\frac{1}{F}\sum_{i=1}^{F}\big{\|}I_{\alpha}^{i}-M_{h}^{i}\big{\|}_{2},$
@@
$L_{reg}$ keeps mouth Gaussians within a predefined cylindrical volume $V$. Let $\{\mathbf{x}_{i}\}_{i=1}^{N}$ be the Gaussian centers inside the mouth. we penalize any point outside $V$ using the signed distance function $SDF(\mathbf{x}_{i},V)$, the equation of $L_{reg}$ loss is as follows:
(6) $L_{reg}=\frac{1}{N}\sum_{i=1}^{N}\left(\max\left(SDF(\mathbf{x}_{i},V),0\right)\right)^{2}$
@@
4. Experiments
4.1. Baseline and Dataset
4.1.1. Baseline
@@
Due to the current lack of methods for synthesizing 3D stylized head avatars using Gaussian blendshapes, we compare our method with the following state‑of‑the‑art methods for synthesizing photo‑realistic 3D head avatars: INSTA (Zielonka et al., 2023), PointAvatar (Zheng et al., 2023), FLARE (Bharadwaj et al., 2023), SplattingAvatar (Shao et al., 2024), FlashAvatar (Xiang et al., 2024), and 3DGB (Ma et al., 2024). Notably, 3DGB shares the similar architecture with ours but focuses on photo‑realistic avatar synthesis, and does not support the synthesis of diverse stylized avatars.
4.1.2. Dataset
@@
We evaluate state-of-the-art methods for photo-realistic avatar synthesis and our method using six videos from INSTA (Zielonka et al., 2023) dataset. Each video is cropped and resized to $512\times 512$ resolution, with sequence lengths ranging from 1,000 to 4,000 frames.
Following the method of 3DGB (Ma et al., 2024), we retain the final 350 frames of each video for testing. Both methods (3DGB (Ma et al., 2024) and ours) apply the same preprocessing techniques (Zielonka et al., 2022, 2023), including background removal and FLAME parameter extraction.
4.2. Evaluation Metrics
@@
We employ two metrics to evaluate video stabilization: Inter-frame Transformation Fidelity (ITF) (Morimoto and Chellappa, 1998; Marcenaro et al., 2001; Xu et al., 2012) and Inter-frame Similarity Index (ISI) (Guilluy et al., 2018; James et al., 2023). ITF measures the inter-frame Peak Signal-to-Noise Ratio (PSNR) in dB based on the mean squared error. The intuitive idea of ITF is that a more stable video (i.e., less jittery) will have greater similarity between adjacent frames compared to an unstable version of the same video. ISI computes the average Structural Similarity (SSIM) between adjacent frames across the video. Higher ISI values indicate greater perceptual similarity between frames, leading to improved visual comfort for viewers.
@@
For 3D head avatar synthesis, we evaluate the performance of state-of-the-art methods for photo-realistic avatar synthesis and our method using standard evaluation metrics (Zhang et al., 2018; Ma et al., 2024), including PSNR (in dB), and SSIM. In addition, we record the training time for each result in minutes and the rendering speed in frames per second (fps).
4.3. Implementation Details
To ensure a fair performance comparison, the training and testing of all methods are performed on a single RTX 4090 GPU. Our methods are implemented in Python using the PyTorch framework. For 2D stylized video generation, we use the pre-trained models provided by StyleGANEX (Yang et al., 2023). For training the 3D stylized head avatars, we employ the Adam optimizer (Kingma and Ba, 2015), setting the initial learning rates of the Gaussian properties $\{\mathbf{x}_{k},\alpha_{k},\mathbf{s}_{k},\mathbf{q}_{k},SH_{k}\}$ to $3.2\times 10^{-7}$, $5\times 10^{-5}$, $5\times 10^{-4}$, $1\times 10^{-4}$, and $1.25\times 10^{-3}$, respectively. Following 3DGB (Ma et al., 2024), the initial number of sampled Gaussians is 50k for the neutral head model and 14k for the mouth interior.

### v2

@@ Abstract
The introduction of 3D Gaussian blendshapes has enabled the real-time reconstruction of animatable head avatars from monocular video. Toonify, a StyleGAN-based method, has become widely used for facial image stylization. To extend Toonify for synthesizing diverse stylized 3D head avatars using Gaussian blendshapes, we propose an efficient two-stage framework, ToonifyGB. In Stage 1 (stylized video generation), we adopt an improved StyleGAN to generate the stylized video from the input video frames, which overcomes the limitation of cropping aligned faces at a fixed resolution as preprocessing for normal StyleGAN. This process provides a more stable stylized video, which enables Gaussian blendshapes to better capture the high-frequency details of the video frames, facilitating the synthesis of high-quality animations in the next stage. In Stage 2 (Gaussian blendshapes synthesis), our method learns a stylized neutral head model and a set of expression blendshapes from the generated stylized video. By combining the neutral head model with expression blendshapes, ToonifyGB can efficiently render stylized avatars with arbitrary expressions. We validate the effectiveness of ToonifyGB on benchmark datasets using two representative styles: Arcane and Pixar.
Introduction
@@
With the advancement of 3D head reconstruction technologies, individuals are now able to personalize unique avatars for telepresence and virtual/augmented reality applications, which serve as an essential foundation for the rise of the metaverse. Considering user preferences and privacy concerns, the creation of stylized avatars is an important topic that deserves further research. Toonify (Pinkney and Adler 2020) is a StyleGAN-based method designed for 2D facial image stylization. In contrast to photo-realistic 3D head avatars, stylized 3D head avatars emphasize the expression of personal identity and the faithful transfer of target styles.
@@
Blendshape is an efficient facial animation representation that synthesizes continuous and high-quality expressions by blending a set of 3D meshes, each corresponding to a specific facial expression. These facial shapes are synthesized by linearly blending the basis meshes using weighting coefficients. With the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al. 2021), Gao et al. (Gao et al. 2022) and Zheng et al. (Zheng et al. 2022) incorporated the blendshape concept into NeRF, enabling avatar animation through the construction of a group of NeRF blendshapes that are linearly blended. Furthermore, the recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) significantly improved rendering efficiency and delivered high-quality head reconstruction, outperforming NeRF in both speed and quality. Based on this, 3D Gaussian Blendshapes (3DGB) (Ma et al. 2024) successfully integrated the blendshape with Gaussian splatting, achieving real-time rendering and state-of-the-art performance in head reconstruction.
@@
In contrast to previous works focused on photo-realistic 3D head avatar reconstruction, we propose ToonifyGB, a two-stage framework for synthesizing and animating 3D stylized head avatars. Given monocular video frames, Stage 1 adopt an improved StyleGAN to generate a more stable and less jittery stylized video, without requiring on fixed resolution and pre-aligned face cropping. In Stage 2, we begin with 3DGB to learn a neutral head model and a set of expression blendshapes, each represented as 3D Gaussians. Finally, by incorporating the facial tracker (Zielonka, Bolkart, and Thies 2022), we use the tracked motion parameters to drive ToonifyGB for animating 3D stylized head avatars.
@@
  • We propose ToonifyGB, an efficient two-stage framework that synthesizes 3D stylized head avatars from monocular videos using Gaussian blendshapes, supporting diverse styles and enabling real-time animation.
@@
  • We demonstrate that reducing per-frame jitter in the generated video enables Gaussian blendshapes to better capture high-frequency details, thereby improving the quality of 3D stylized head avatar animations.
@@
  • To the best of our knowledge, this work is the first to synthesize 3D stylized head avatars based on Gaussian blendshapes.
@@
Figure 1: Pipeline: Our ToonifyGB framework consists of two stages: Stage 1 involves the generation of stylized videos, and Stage 2 focuses on the synthesis of 3D stylized head avatars using Gaussian blendshapes.
Related Work
StyleGAN and Toonify
StyleGAN (Karras, Laine, and Aila 2019; Karras et al. 2020) has been widely used to generate realistic facial images across diverse styles. Inversion of StyleGAN enables projecting real facial images into its latent space, allowing for subsequent edits such as adding glasses or changing hairstyle or age (Abdal, Qin, and Wonka 2019; Patashnik et al. 2021). To enhance inversion efficiency, methods such as pSp (Richardson et al. 2021) and e4e (Tov et al. 2021) employ encoders to directly project target faces into their corresponding latent codes. However, these methods often struggle to reconstruct fine image details, resulting in unsatisfactory reconstruction quality. To address these limitations, ReStyle (Alaluf, Patashnik, and Cohen-Or 2021) and HFGI (Wang et al. 2022) improve reconstruction fidelity by respectively predicting latent code residuals and correcting middle layer features. Nevertheless, these methods remain limited to aligned and cropped facial images to achieve effective editing and reconstruction.
Recently, researchers (Pinkney and Adler 2020; Ojha et al. 2021; Jang et al. 2021; Yang et al. 2022a; Gal et al. 2022) have explored using StyleGAN for target‑domain image generation through transfer learning. Among these works, Toonify (Pinkney and Adler 2020) fine‑tunes the trained generator to blend realistic textures with toonified facial structures. In addition to image editing, StyleGAN has also been widely applied to video editing. Related studies have focused on enhancing video editing performance by employing temporal correlations in low‑dimensional latent codes (Fox et al. 2021), disentangling identity from facial attributes (Yao et al. 2021), incorporating sketch‑based branches (Liu et al. 2022), and tuning the generator to maintain temporal consistency (Tzaban et al. 2022). However, these methods typically rely on face alignment and cropping as preprocessing. Although StyleGAN3 (Karras et al. 2021) was introduced to support unaligned face inputs, a subsequent study (Alaluf et al. 2022) has shown that it struggles to encode facial features effectively without preprocessing, often resulting in structural artifacts. To overcome these limitations, methods such as VToonify (Yang et al. 2022b) and StyleGANEX (Yang et al. 2023) have been proposed to directly process videos beyond pre-aligned cropping. Nevertheless, these methods remain limited to 2D representations and have yet to be extended to 3D applications.
3D Head Avatar
@@
Since the introduction of NeRF (Mildenhall et al. 2021), implicit representation-based methods (Yenamandra et al. 2021; Zheng et al. 2022; Hong et al. 2022; Chan et al. 2022; Xu et al. 2023) for head reconstruction have achieved remarkable progress. 3DGS (Kerbl et al. 2023) has obtained a significant breakthrough in 3D reconstruction, further advancing the development of downstream applications such as 3D head modeling. Although several Gaussian-based methods (Qian et al. 2024; Xu et al. 2024; Chen et al. 2024; Ma et al. 2024; Xiang et al. 2024; Abdal et al. 2024; Kirschstein et al. 2024) have demonstrated high-quality head reconstruction and impressive rendering performance, they typically focus on photo-realistic avatars, with relatively limited exploration of avatar stylization. Stylized head avatars, characterized by geometric abstraction and artistic expression, differ significantly from the photo-realistic avatars synthesized by the aforementioned methods.
@@
Pre-trained 3D GANs (Wu et al. 2016) enable high-quality generation, making 3D head stylization possible. Although fine-tuning 3D generators for geometric and texture-based stylization has proven effective (Or-El et al. 2022; Jin et al. 2022; Abdal et al. 2023; Lan et al. 2023; Wang et al. 2023; Zhang et al. 2024), performing independent fine-tuning for each new style remains costly. Toonify3D (Jang et al. 2024) addressed this limitation by predicting facial surface normals using the proposed StyleNormal, enabling direct face stylization without additional fine-tuning. Similarly, DeformToon3D (Zhang et al. 2023) introduced StyleField to predict conditional 3D deformations, aligning NeRF representations in real space with style space to achieve geometric stylization and obviate per-style fine‑tuning. However, Toonify3D suffers from limited data diversity, and DeformToon3D cannot support novel-view animations, which limits their application scenarios.
Method
ToonifyGB Framework
@@
Given a monocular video input, ToonifyGB applies frame-by-frame stylization to generate the corresponding stylized frames. For inputs such as live streams or selfie videos, the face often occupies only a small portion of each frame, while the rest includes the hairstyle and upper body. Traditional methods (Karras, Laine, and Aila 2019; Karras et al. 2020) typically require face alignment, cropping, and editing before synthesizing the results back into the original frame. This process often introduces visual discontinuities at the seams, resulting in noticeable jitter in the output video. To address this issue, we adopt an improved StyleGAN model based on StyleGANEX (Yang et al. 2023) in Stage 1, enabling stable stylized video generation at the original resolution, as shown in Figure 1.
@@
To prepare the training data for Stage 2, we follow the method in (Zielonka, Bolkart, and Thies 2023; Ma et al. 2024), using the facial tracker from (Zielonka, Bolkart, and Thies 2022) to compute FLAME (Li et al. 2017) meshes, including a neutral head model and a set of expression blendshapes. This process also provides camera parameters $C$, joint and pose parameters $\Theta$, and expression coefficients $\{\psi_{k}\}$ for each frame. In addition to enabling facial expressions control, the FLAME model based on Principal Component Analysis (PCA) provides joint and pose parameters for controlling head, eyeball, eyelid, and jaw movements. As shown in Figure 1, we apply Linear Blend Skinning (LBS) to transform the Gaussian model based on the extracted joint and pose parameters. The transformation is defined as:
$B^{\psi*}=LBS(B^{\psi},\Theta).$ (1)
The transformed Gaussian model is then rendered in real-time as a 3D stylized head avatar using Gaussian Splatting. Finally, by integrating the camera parameters $C$, we enable novel-view rendering and animation.
Figure 2: Visualization of stylized video generation results on the INSTA (Zielonka, Bolkart, and Thies 2023) and NeRFBlendShape (Gao et al. 2022) datasets.
Stylized Video Generation
As shown in StyleGANEX (Yang et al. 2023), manipulating feature maps at different layers of StyleGAN leads to different spatial effects in the generated faces. Specifically, while shifting or rotating the feature maps in deeper layers (i.e., Layer 7) produces consistent global transformations, similar operations in shallow layers (i.e., Layer 1) fail to preserve facial structure due to the low spatial resolution of the $4\times 4$ feature map, causing blurring and loss of detail. To address this limitation, we adopt StyleGANEX (Yang et al. 2023), an enhanced variant of StyleGAN2 (Karras, Laine, and Aila 2019), which increases the spatial resolution of shallow feature maps (Layers 1–7) to $32\times 32$. This improvement enables finer control over facial geometry and enhances the generation quality for unaligned faces.
@@
Our specific architectural improvements of the generator are as follows. First, we replace the constant $4\times 4$ input of the first layer with a variable feature map of resolution $1/32$ of the final output, enabling support for arbitrary input sizes. Then, we replace the standard convolutions in the shallow layers with dilated convolutions to enlarge the receptive field. Finally, we remove all upsample operations before the eighth layer, ensuring that the seven shallow layers maintain the same $32\times 32$ resolution.
These architectural improvements effectively address the limitations beyond pre-aligned cropping. As shown in Figure 2, our method consistently generates high-quality stylized head videos across diverse styles, regardless of gender.
Gaussian Blendshapes Synthesis
@@
We represent all Gaussian head avatars using 3D Gaussians. As described in 3DGS (Kerbl et al. 2023), each Gaussian has some basic properties including Gaussian center $\mu$, scale $s$, color $c$, opacity $\alpha$, and rotation $q$. Based on 3DGB (Ma et al. 2024), our Gaussian blendshape representation consists of a neutral model $B_{0}$ and a set of $n$ expression blendshapes ${B_{1},B_{2},\ldots,B_{n}}$. Each Gaussian in the neutral model $B_{0}$ has a set of blend weights $w$ to control joint and pose. In addition, each Gaussian in an expression blendshape $B_{k}$ corresponds one‑to‑one to a Gaussian in the neutral model $B_{0}$. The difference between $B_{k}$ and $B_{0}$ is defined as the difference in their corresponding Gaussian properties: $\Delta B_{k}=B_{k}-B_{0}$. The expression of Gaussian head avatar $B^{\psi}$ can be computed as follows:
$B^{\psi}=B_{0}+\sum_{k=1}^{n}\psi_{k}\,\Delta B_{k}$ (2)
where $\psi_{k}$ denotes the expression coefficients. Here, $B^{\psi}$ represents the untransformed expression model, and the final posed Gaussian model, obtained via Linear Blend Skinning (LBS), is defined in Equation 1.
@@
Since the FLAME meshes and blendshape models do not include interior mouth components such as teeth, we adopt the method of 3DGB (Ma et al. 2024) by defining a separate set of Gaussians for the mouth $B_{m}$. The properties of these mouth Gaussians are not affected by expression changes, they only move with the jaw joint in the FLAME model. The mouth Gaussians for the head avatar, $B_{m}^{*}$, are computed via linear blend skinning (LBS) as:
$B_{m}^{*}=LBS(B_{m},\Theta).$ (3)
The transformed Gaussian model $(B^{\psi*},B_{m}^{*})$ is rendered into a complete 3D head avatar using real-time Gaussian Splatting, with the overall pipeline shown in Figure 1.
Loss Function
We adopt the loss function from 3DGB (Ma et al. 2024), and define the total loss as follows:
$L=\lambda_{1}L_{rgb}+\lambda_{2}L_{\alpha}+\lambda_{3}L_{reg},$ (4)
where the default weights of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are set to 1, 10, 100, respectively.
@@
The RGB loss $L_{rgb}$ encourages the rendered image to resemble the target video frame in both color and structure. It is computed as a weighted combination of an $L_{1}$ loss and a differentiable Structural Similarity (D-SSIM) loss:
$L_{rgb}=\lambda_{rgb}L_{1}+(1-\lambda_{rgb})L_{D-SSIM},$ (5)
where the default weight $\lambda_{rgb}$ is set to $0.2$.
@@
The opacity loss $L_{\alpha}$ penalizes opacity values outside the head mask. For each frame $i$, we compute the accumulated opacity image $I_{\alpha}^{i}$ and the corresponding head mask $M_{h}^{i}$, and average the error over $F$ frames:
$L_{\alpha}=\frac{1}{F}\sum_{i=1}^{F}\left\|I_{\alpha}^{i}-M_{h}^{i}\right\|_{2}.$ (6)
@@
The regularization loss $L_{reg}$ constrains the mouth Gaussians to remain within a predefined cylindrical volume $V$. Let $\{\mathbf{x}_{i}\}_{i=1}^{N}$ denote the centers of Gaussians located in the mouth region. To penalize points outside the volume, we employ a signed distance function $SDF(\mathbf{x}_{i},V)$, and define the loss as follows:
$L_{reg}=\frac{1}{N}\sum_{i=1}^{N}\left(\max\left(SDF(\mathbf{x}_{i},V),0\right)\right)^{2},$ (7)
@@
Experiments
Baseline and Dataset
@@
Due to the current lack of methods for synthesizing 3D stylized head avatars using Gaussian blendshapes, we compare our method against the following state-of-the-art methods for photo-realistic 3D head avatar synthesis: INSTA (Zielonka, Bolkart, and Thies 2023), PointAvatar (Zheng et al. 2023), FLARE (Bharadwaj et al. 2023), SplattingAvatar (Shao et al. 2024), FlashAvatar (Xiang et al. 2024), and 3DGB (Ma et al. 2024). Notably, 3DGB shares a similar architecture with ours but focuses on photo-realistic avatar synthesis and does not support the synthesis of diverse stylized avatars.
@@
We evaluate both our method and state-of-the-art photo-realistic avatar synthesis methods using six videos from the INSTA (Zielonka, Bolkart, and Thies 2023) dataset. Each video is cropped and resized to $512\times 512$ resolution, with sequence lengths ranging from 1,000 to 4,000 frames. Following the method of 3DGB (Ma et al. 2024), we retain the final 350 frames of each video for testing. Both 3DGB (Ma et al. 2024) and our method apply the same preprocessing pipeline (Zielonka, Bolkart, and Thies 2022, 2023), including background removal and FLAME parameter extraction.
Evaluation Metrics
@@
We employ two metrics to evaluate video stabilization: Inter-frame Transformation Fidelity (ITF) (Morimoto and Chellappa 1998; Marcenaro, Vernazza, and Regazzoni 2001; Xu et al. 2012) and Inter-frame Similarity Index (ISI) (Guilluy, Beghdadi, and Oudre 2018; James, Jain, and Rajwade 2023). ITF measures the inter-frame Peak Signal-to-Noise Ratio (PSNR) in dB based on the mean squared error. The intuitive idea of ITF is that a more stable video (i.e., less jittery) will have greater similarity between adjacent frames compared to an unstable version of the same video. ISI computes the average Structural Similarity (SSIM) between adjacent frames across the video. Higher ISI values indicate greater perceptual similarity between frames, leading to improved visual comfort for viewers.
@@
For 3D head avatar synthesis, we evaluate the performance of our method and state-of-the-art methods for photo-realistic avatar synthesis using standard evaluation metrics (Zhang et al. 2018; Ma et al. 2024), including Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM). In addition, we record the training time (in minutes) and the rendering speed (in frames per second, fps) for each method. In the ablation study, we additionally adopt the Learned Perceptual Image Patch Similarity (LPIPS) metric to better capture perceptual differences between the synthesized avatars and the ground truth.